A car begins at rest and accelerates up to 20.3 m/s in 10 seconds. how far did it travel in this time?


1
Expert's answer
2021-02-17T15:23:00-0500
"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\cdot(0+20.3\ \dfrac{m}{s})\cdot10\ s=101.5\ m."